If Spain specializes in production it can produce 300 Tons of Fish or 327 Tons of Grain. Norway on the other hand can produce 200 Tons of Fish or 200 Tons of Grain. Calculate the ratio of Tons of Fish relative to Tons of Grain associated with the country with the least opportunity cost to produce grain. Record your answer to two decimal places.
Answer #1

Opportunity cost of Producing Grain for

Spain = 300/327 = 0.92 ton of fish

Norway = 200/200 =1 ton of fish

So, Spain has the least opportunity cost to produce grain = 0.92 ton of fish per ton of grain